OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 i think there is no solution

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

2x + 6 = 24 2x + 6 = -24

OpenStudy (anonymous):

solve for x

OpenStudy (abmon98):

Isolate absolute \[\left| 2x+6 \right|=24\] \[\left| ax+c \right|=b \rightarrow ax+c=-b \rightarrow ax+c=b\] \[2x+6=24\]\[2x+6=-24\] \[2x=18\]\[2x=-30\] \[x=9\]\[x=-15\]